# Two Optimizers Are Better Than One: LLM Catalyst Empowers Gradient-Based Optimization for Prompt Tuning

### v3

@@ 3.3 Catalyzing Prompt Tuning by Using LLM as a Prompt Optimizer
Our method optimizes the prompt by using the gradient optimizer and MaaO in an alternating pattern. Specifically, we first update the parameter of the prompt for minor steps of gradient-descent optimization and record the intermediate learned prompts and corresponding fitness scores, which are evaluated on the few-shot training samples. Then, we construct instruction for LLM with the intermediate learned prompts as optimizing trajectory information. Taking the instruction as input, LLM generates more promising candidate prompts for the target model. Next, we reinitialize the parameter of the prompts with LLM-generated prompts and restart the gradient-based training process for the next round. After operating the above two optimizers alternately for few rounds, we finally train the prompt to convergence using the gradient-based optimizer. In the following, we will describe the components of MaaO and show the combination of it with the gradient-based optimizer for a collaborative prompt optimization.
@@
Instruction construction. The gradient optimizer calculates updates based on the current parameters and objective function. Information of the current state of optimization should also be properly provided for LLM to infer from. We collect the intermediate optimized prompt in the training trajectory of the gradient-based optimizer and evaluate the performance corresponding to each intermediate prompt as a fitness score, indicating how good or bad the prompt performs. Considering that the accuracy may be not precise enough on few samples, we employ loss as the indicator value. LLM is instructed to generate prompts that potentially achieve better performance based on observed patterns in top-performing candidates.
@@
We also briefly define the role of LLM and explain the goal of optimization in natural language, encouraging LLM to assemble task-related tokens when constructing the prompt. Additional instructions to constrain the length and number of the generated prompts are included for programmed processing. Figure 3 shows the instruction used in each optimization iteration of prompt using GPT-3.5 and GPT-4.
@@
Token space projection. Gradient-based prompt tuning typically optimizes soft prompt tokens in the embedding space of the pre-trained model. However, it is not feasible to directly provide the soft embedding vectors as input to the LLM, which is proficient in responding to natural language with semantics. To convert the soft prompt embedding to discrete words, we employ a reverse process of word2vec (Mikolov et al., 2013) to project the embedding to the matched vocabulary.
@@
Given a pre-trained target model with token embedding layer $\mathcal{V\left(\cdot\right)}$, textual inputs $\{t_{i}\}_{i=1}^{l}$ to the model are first converted to vector sequence as $\{\bm{t}_{i}\}_{i=1}^{l}=\{\mathcal{V}(t_{i})|i\in[1,l]\}$, before input into the model. Gradient-based prompt tuning optimizes in the continuous vector space for best prompt embeddings. We project the continuous prompt vector $\bm{\hat{t}}_{i}$ to nearest discrete tokens by $\hat{t_{i}}=\mathcal{V}^{-1}(\bm{\hat{t}}_{i})$, where $\mathcal{V}^{-1}$ is defined as:
@@
$\mathcal{S}$ denotes the dictionary of the pre-trained model. The projected prompt is used to construct the instruction for LLM to infer for better prompt candidates in the unconstrained semantic vocabulary space.
Integration with gradient-based optimizer. Gradient-based optimizer conduct rigorous local-optimal update on the parameter based by back-propagated gradient. MaaO infers for promising candidate prompts by analyzing and generating semantic-related prompts based on currently found solutions. We propose a cooperation workflow of the gradient-based optimizer and MaaO in Algorithm 1.
We connect the two optimizers in two ways. First, the gradient optimizer provides the LLM with the intermediate results in the prompt optimization process, from which LLMs infer for more promising candidate prompts. The generated prompts by LLMs assemble task-related semantic contents and provide opportunities to break free from local optimal that may encountered in gradient-based optimization. Second, we restart the gradient-based optimization by using the prompts generated by the LLM optimizer as new initial values of the gradient optimizer to obtain refined prompts based on the LLM-generated ones. By optimizing the prompt based on the two optimizers alternately, it guides the LLM to progressively exploit better prompts in a more promising area of the search space near the previously found good solutions. The gradient optimizer provides stable convergence for the final learned prompts. Note that the overhead brought by our optimization algorithm compared to the original gradient-based prompt tuning is only about dozens (at most 30) of iterations using combined optimizer.
@@
Table 1: Few-shot prompt optimization results on downstream datasets.
Datasets Settings ResNet50 ViT-B/16
CoOp BBO Ours(GPT-3.5) Ours(GPT-4) Ours(Llama7B) TCP Ours(GPT-3.5)
Eurosat 1-shot 50.58 49.0 56.27 56.74 55.38 65.04 63.06
4-shot 69.65 - 71.17 72.55 73.39 72.42 77.35
8-shot 72.74 - 74.33 75.99 76.79 77.71 79.49
16-shot 83.57 51.4 83.77 85.07 83.95 84.43 86.18
Avg. 69.14 71.39 72.59 72.38 74.90 76.52
DTD 1-shot 43.13 44.8 47.24 44.78 42.47 55.06 55.14
4-shot 53.45 - 54.87 55.04 54.14 61.88 63.36
8-shot 59.38 - 60.30 60.18 60.48 68.62 68.56
16-shot 63.87 44.9 64.40 64.30 64.48 73.48 73.66
Avg. 54.96 - 56.70 56.08 55.39 64.76 65.18
Caltech101 1-shot 87.76 89.1 87.02 86.87 87.86 94.08 94.00
4-shot 89.05 - 88.72 88.68 89.03 95.18 95.38
8-shot 90.58 - 90.26 90.66 90.25 95.39 95.33
16-shot 91.66 89.5 92.28 91.83 92.33 95.89 95.71
Avg. 89.76 - 89.57 89.51 89.87 95.14 95.11
Oxford_Flowers 1-shot 69.09 67.2 71.62 71.28 72.31 85.80 87.05
4-shot 87.00 - 89.38 88.51 88.93 94.72 95.09
8-shot 90.19 - 91.15 90.77 90.41 96.14 96.31
16-shot 93.88 67.4 94.42 94.02 94.49 97.47 97.54
Avg. 85.04 - 86.64 86.15 86.54 93.53 94.00
Fgvc_Aircraft 1-shot 18.38 18.1 18.69 18.82 18.50 28.90 28.33
4-shot 21.90 - 22.73 22.77 23.19 35.61 35.61
8-shot 25.15 - 26.53 26.42 27.35 39.76 40.69
16-shot 28.86 18.1 31.27 31.24 31.44 43.29 43.79
Avg. 23.57 - 24.81 24.81 25.12 36.89 37.11
Food101 1-shot 72.60 78.3 73.86 73.98 72.47 85.74 85.40
4-shot 70.93 - 70.44 70.25 69.76 86.43 86.01
8-shot 73.97 - 73.12 73.97 72.51 86.83 86.67
16-shot 75.72 78.3 75.20 74.94 73.85 87.25 87.26
Avg. 73.31 - 73.16 73.29 72.15 86.56 86.34
Stanford_Cars 1-shot 55.70 56.2 54.74 54.89 54.78 68.87 68.05
4-shot 61.22 - 61.93 61.76 62.01 75.25 76.17
8-shot 65.14 - 65.89 66.82 67.37 79.27 79.29
16-shot 67.97 56.8 68.84 69.34 73.30 83.79 83.98
Avg. 62.51 - 62.85 63.20 64.37 76.80 76.87
Oxford_Pets 1-shot 85.14 88.1 85.70 84.38 84.78 91.26 90.75
4-shot 85.37 - 85.03 85.33 84.47 92.67 92.65
8-shot 85.70 - 84.65 84.82 84.58 92.91 92.55
16-shot 86.84 88.3 86.37 86.00 85.02 93.34 93.19
Avg. 85.76 - 85.44 85.13 84.71 92.55 92.29
UCF101 1-shot 62.60 60.2 61.85 61.07 62.34 73.44 72.69
4-shot 68.75 - 68.25 69.18 68.27 80.93 80.77
8-shot 72.26 - 72.69 72.26 72.58 83.18 83.40
16-shot 74.91 60.5 74.82 75.96 74.95 85.25 85.13
Avg. 69.63 - 69.40 69.62 69.54 80.70 80.50
SUN397 1-shot 58.33 61.0 58.13 57.25 57.47 69.20 69.13
4-shot 64.48 - 65.21 64.06 64.40 73.78 73.94
8-shot 66.79 - 67.20 66.90 66.79 75.78 75.99
16-shot 68.79 60.8 68.39 68.42 68.42 76.81 76.69
Avg. 64.60 - 64.73 64.16 64.27 73.89 73.94
@@ 4.1 Experimental Setup
Implementation details and baselines. To comprehensively evaluate the effectiveness of our method, we use GPT-3.5, GPT-4 (OpenAI et al., 2024), and white-box LLM Llama2 (Touvron et al., 2023b) as the LLM-based optimizer. We integrate MaaO into existing prompt tuning methods for language pre-trained models, P-tuning (Liu et al., 2023b),  Lester et al. (2021), and two methods for vision-language model, CoOp (Zhou et al., 2022b), TCP (Yao et al., 2024). CoOp pioneered prompt tuning for vision-language models, and TCP is one of the recent state-of-the-art prompt tuning approaches. Both of the methods are typical textual prompting methods for vision-language models. For a fair comparison, we fix the original hyperparameter of previous methods, such as pre-trained backbone and prompt module design, and only apply our method as new optimization strategy. For the configuration of Algorithm 1, the number of rounds $N$ is set as 3 and the iteration for the gradient optimizer $m$ is set as 10. All experimental results are averaged over 3 random seeds. The experiments are conducted on a V100 GPU. Detailed hyperparameter setting is provided in the Appendix.
Datasets. For lauguage model, we conduct experiments over commonly-used pre-trained model RoBERTa (Liu et al., 2019) on NLU tasks from SuperGLUE (Wang et al., 2020) to test our methods. We apply our prompt optimization algorithm to vision-language pretrained CLIP (Radford et al., 2021) for adaptation of image classification tasks. We adopt commonly used 10 datasets to comprehensively evaluate our method, including Caltech101 (Li et al., 2004), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback and Zisserman, 2008), Food101 (Bossard et al., 2014), FGVCAircraft (Maji et al., 2013), SUN397 (Xiao et al., 2010), UCF101 (Soomro et al., 2012), DTD (Cimpoi et al., 2014), and EuroSAT (Helber et al., 2019). For each dataset, labeled few-shot samples from each class are used as training data.
@@ 4.2 Main Results
Experiments on vision-language pre-trained models. We compare the results of employing our proposed optimization method for existing prompting methods (Zhou et al., 2022b; Yao et al., 2024). From Table 1, the results on "RN50" backbone show that our integrated optimization outperforms existing gradient-based prompt tuning methods at seven out of ten benchmark datasets and the other tasks remain close to the baseline performance. Both black-box GPT and open-sourced Llama2 achieve consistent improvements, demonstrating the effectiveness of our combined optimization framework. TCP (Yao et al., 2024) is one of the state-of-the-art prompt tuning approaches with a stronger backbone. Although the absolute improvement inevitably decreases, our method still brings stable improvements on six out of ten datasets.
We also compare with methods that optimize by LLM only, e.g., BBO (Liu et al., 2023a). We list the results reported in BBO paper since the code has not been released yet and our reproduced results can not match the results published in the paper. Although BBO achieves a completely program-free prompt learning method for CLIP, their performances in 16-shot setting are poor. Their method largely relies on the inherent deductive ability of LLM, which can not make good use of information in more training samples. Our method instruct LLMs with intermediate results of gradient optimizer, enabling LLMs to exploit in a more promising sub-region of the semantic space. The prompts generated by LLM are further refined by the disciplined gradient-based optimizer to achieve stable improvements.
Experiments on language models. We also test our method on pre-trained language models and NLU tasks. From Table 3, our method with GPT-4 as optimizer surpasses P-tuning (Liu et al., 2023b) and  Lester et al. (2021) on four out of five tasks from SuperGLUE, showing the superiority of our optimization method over vanilla gradient-based prompt tuning.
@@
Table 2: Results on SuperGLUE dev-set. (PT: P-tuning & Lester et al. (2021); Model: RoBERTa-Large).
Methods COPA BoolQ RTE WiC WSC
PT 61.67 62.29 55.72 53.81 64.10
Ours 68.67 63.09 58.00 55.85 63.46

### v4

@@ 3.3 A Collaborative Optimization Method by Using LLM as a Prompt Optimizer
Our method optimizes the prompt by using the gradient optimizer and MaaO in an alternating pattern. Specifically, we first update the parameter of the prompt for minor steps of gradient-descent optimization and record the intermediate learned prompts and corresponding fitness scores, which are evaluated on the few-shot training samples. Then, we construct instruction for LLM with the intermediate learned prompts as optimizing trajectory information. Taking the instruction as input, LLM generates more promising candidate prompts for the target model. Next, we reinitialize the parameter of the prompts with LLM-generated prompts and restart the gradient-based training process for the next round. After operating the above two optimizers alternately for few rounds, we finally train the prompt to convergence using the gradient-based optimizer. In the following, we will describe the components of MaaO and show their combination with the gradient-based optimizer for collaborative prompt optimization.
@@
Instruction construction. The gradient optimizer calculates updates based on the current parameters and objective function. Information on the current state of optimization should also be properly provided for LLM to infer from. We collect the intermediate optimized prompt in the training trajectory of the gradient-based optimizer and evaluate the performance corresponding to each intermediate prompt as a fitness score, indicating how good or bad the prompt performs. Considering that the accuracy may not be precise enough on a few samples, we employ loss as the indicator value. LLM is instructed to generate prompts that potentially achieve better performance based on observed patterns in top-performing candidates.
@@
We also briefly define the role of LLM and explain the optimization goal in natural language, encouraging LLM to assemble task-related tokens when constructing the prompt. Additional instructions to constrain the length and number of the generated prompts are included for programmed processing. Figure 3 shows the instruction used in each optimization iteration of the prompt using GPT-3.5 and GPT-4.
@@
Token space projection. Gradient-based prompt tuning typically optimizes continuous prompt embeddings in the token space of the pre-trained model. However, it is not feasible to directly provide the soft embedding vectors as input to the LLM, which is proficient in responding to natural language with semantics. To convert the soft prompt embedding to discrete words, we employ a reverse process of word2vec (Mikolov et al., 2013) to project the embedding to the matched vocabulary.
@@
Given a pre-trained target model with token embedding layer $\mathcal{V\left(\cdot\right)}$, textual inputs $\{t_{i}\}_{i=1}^{l}$ to the model are first converted to vector sequence as $\{\bm{t}_{i}\}_{i=1}^{l}=\{\mathcal{V}(t_{i})|i\in[1,l]\}$, before input into the model. Gradient-based prompt tuning optimizes in the continuous vector space for best prompt embeddings. We define an inverse projection function $\mathcal{V}^{-1}(\cdot)$ to project the continuous prompt vector $\bm{\hat{t}}_{i}$ to nearest discrete tokens by $\hat{t_{i}}=\mathcal{V}^{-1}(\bm{\hat{t}}_{i})$. $\mathcal{V}^{-1}$ is defined as:
@@
$\mathcal{S}$ denotes the dictionary of the pre-trained model. The projected prompt is used to construct the instruction for LLM to infer better prompt candidates in the unconstrained semantic vocabulary space.
Integration with gradient-based optimizer. Gradient-based optimizers conduct rigorous local-optimal updates on the parameters based on back-propagated gradient. MaaO infers promising candidate prompts by analyzing and generating semantic-related prompts based on currently found solutions. We propose a cooperation workflow of the gradient-based optimizer and MaaO in Algorithm 1.
We connect the two optimizers in two ways. First, the gradient optimizer provides the LLM with the intermediate results in the prompt optimization process, from which LLMs infer more promising candidate prompts. The generated prompts by LLMs assemble task-related semantic contents and provide opportunities to break free from local optimal that may encountered in gradient-based optimization. Second, we restart the gradient-based optimization by using the prompts generated by the LLM optimizer as new initial values of the gradient optimizer to obtain refined prompts based on the LLM-generated ones. Optimizing the prompt based on the two optimizers alternately guides the LLM to progressively exploit better prompts in a more promising area of the search space near the previously found good solutions. The gradient optimizer provides stable convergence for the final learned prompts. Note that the overhead brought by our optimization algorithm compared to the original gradient-based prompt tuning is only about dozens (at most 30) of iterations using the combined optimizer.
@@
Table 1: Results of prompt tuning pre-trained language model RoBERTa-Large on SuperGLUE dev-set. (PT: P-tuning & Lester et al. (2021)).
Methods COPA BoolQ RTE WiC WSC Avg.
PT 61.67 62.29 55.72 53.81 64.10 59.52
Ours 68.67 63.09 58.00 55.85 63.46 61.81
Table 2: Result of few-shot prompt tuning vision-language model CLIP on downstream datasets. Top optimization results of different optimizer LLMs are marked with different colors.
Datasets Settings ResNet50 ResNet50 ViT-B/16
CoOp Liu et al.(2023a) Ours(GPT3.5) Ours(GPT4) Ours(Llama) TCP Ours(GPT3.5) TCP Ours(GPT3.5)
Eurosat 1-shot 50.582.74 49.0 56.272.06 56.742.28 55.381.91 62.792.10 63.070.56 65.040.99 63.061.05
4-shot 69.650.73 - 71.170.84 72.550.87 73.390.89 73.201.40 74.100.39 72.420.50 77.350.26
8-shot 72.740.96 - 74.331.90 75.991.05 76.790.67 77.370.44 77.950.15 77.710.02 79.490.19
16-shot 83.570.46 51.4 83.771.18 85.070.55 83.950.58 82.370.29 83.010.41 84.430.09 86.180.15
Avg. 69.14 - 71.39 72.59 72.38 73.93 74.53 74.90 76.52
DTD 1-shot 43.131.86 44.8 47.240.37 44.781.29 42.471.09 48.250.32 48.640.25 55.061.20 55.140.35
4-shot 53.450.47 - 54.870.90 55.040.71 54.140.21 60.280.24 60.300.27 61.880.05 63.360.44
8-shot 59.381.06 - 60.300.76 60.180.29 60.480.61 64.380.53 65.260.20 68.620.46 68.560.27
16-shot 63.870.24 44.9 64.400.23 64.300.88 64.480.73 68.180.61 68.450.26 73.480.14 73.660.12
Avg. 54.96 - 56.70 56.08 55.39 60.27 60.66 64.76 65.18
Caltech101 1-shot 87.760.92 89.1 87.020.56 86.87 0.58 87.860.39 89.160.43 89.580.15 94.080.23 94.000.18
4-shot 89.050.55 - 88.720.27 88.680.24 89.030.16 91.150.05 91.400.23 95.180.02 95.380.17
8-shot 90.580.52 - 90.260.59 90.660.41 90.250.94 92.050.26 91.990.20 95.390.19 95.330.07
16-shot 91.660.22 89.5 92.280.50 91.830.05 92.330.06 93.250.20 93.050.02 95.890.16 95.710.15
Avg. 89.76 - 89.57 89.51 89.87 91.40 91.51 95.14 95.11
Oxford Flowers 1-shot 69.091.57 67.2 71.621.01 71.281.03 72.311.22 78.510.37 77.940.69 85.800.56 87.050.36
4-shot 87.000.91 - 89.380.75 88.510.43 88.930.52 90.850.13 91.030.20 94.720.20 95.090.28
8-shot 90.190.34 - 91.150.63 90.770.45 90.410.39 93.310.05 93.740.18 96.140.14 96.310.04
16-shot 93.880.13 67.4 94.420.28 94.020.31 94.490.41 95.380.21 95.330.15 97.470.05 97.540.10
Avg. 85.04 - 86.64 86.15 86.54 89.51 89.51 93.53 94.00
Fgvc Aircraft 1-shot 18.380.84 18.1 18.690.67 18.820.48 18.500.38 20.010.28 20.690.14 28.900.31 28.330.21
4-shot 21.900.57 - 22.731.07 22.770.34 23.191.03 25.180.16 25.720.11 35.610.60 35.610.31
8-shot 25.150.55 - 26.530.16 26.420.54 27.350.56 29.970.24 30.310.18 39.760.49 40.690.32
16-shot 28.860.59 18.1 31.270.12 31.240.56 31.440.78 34.030.69 34.360.14 43.290.22 43.790.29
Avg. 23.57 - 24.81 24.81 25.12 27.30 27.77 36.89 37.11
Food101 1-shot 72.600.75 78.3 73.860.40 73.980.53 72.470.70 75.900.20 75.480.13 85.740.12 85.400.16
4-shot 70.930.41 - 70.440.36 70.250.55 69.7640 76.090.13 75.990.21 86.430.15 86.010.08
8-shot 73.970.51 - 73.120.13 73.970.30 72.510.11 77.340.12 77.110.17 86.830.01 86.670.11
16-shot 75.720.15 78.3 75.200.26 74.940.19 73.850.21 78.470.09 78.440.02 87.250.15 87.260.06
Avg. 73.31 - 73.16 73.29 72.15 76.95 76.76 86.56 86.34
Stanford Cars 1-shot 55.700.56 56.2 54.740.81 54.890.93 54.780.62 56.370.28 55.590.27 68.870.79 68.050.67
4-shot 61.220.54 - 61.930.18 61.760.17 62.010.42 66.020.26 66.870.16 75.250.26 76.170.20
8-shot 65.140.54 - 65.890.42 66.820.59 67.370.60 71.020.30 70.800.37 79.270.29 79.290.21
16-shot 67.970.36 56.8 68.841.10 69.340.38 73.300.28 75.390.20 75.810.13 83.790.11 83.980.23
Avg. 62.51 - 62.85 63.20 64.37 67.20 67.27 76.80 76.87
Oxford Pets 1-shot 85.140.78 88.1 85.700.93 84.380.36 84.780.53 86.760.31 86.160.11 91.260.40 90.750.24
4-shot 85.370.44 - 85.030.86 85.330.88 84.470.67 88.290.17 87.470.03 92.670.25 92.650.11
8-shot 85.700.53 - 84.650.20 84.820.39 84.580.21 87.850.10 87.440.13 92.910.27 92.550.13
16-shot 86.840.08 88.3 86.370.07 86.000.32 85.020.24 89.630.26 89.230.13 93.340.07 93.190.18
Avg. 85.76 - 85.44 85.13 84.71 88.13 87.58 92.55 92.29
UCF101 1-shot 62.600.88 60.2 61.850.69 62.120.09 62.340.45 64.720.74 64.510.44 73.440.54 72.690.11
4-shot 68.750.38 - 68.250.56 69.180.79 68.270.84 72.570.51 73.850.11 80.930.08 80.770.09
8-shot 72.260.30 - 72.690.19 72.260.44 72.580.41 76.680.01 77.610.19 83.180.20 83.400.10
16-shot 74.910.33 60.5 74.820.87 75.960.14 74.950.33 78.910.20 79.950.01 85.250.25 85.130.27
Avg. 69.63 - 69.40 69.88 69.54 73.22 73.98 80.70 80.50
SUN397 1-shot 58.330.76 61.0 58.130.65 57.250.43 57.470.79 60.940.17 61.130.22 69.200.19 69.130.12
4-shot 64.480.25 - 65.210.37 64.060.40 64.400.39 67.120.04 67.370.11 73.780.04 73.940.10
8-shot 66.790.23 - 67.200.24 66.900.22 66.790.33 69.830.12 69.740.09 75.780.01 75.990.04
16-shot 68.790.26 60.8 68.390.09 68.420.18 68.420.21 71.970.18 72.260.15 76.810.12 76.690.07
Avg. 64.60 - 64.73 64.16 64.27 67.47 67.63 73.89 73.94
@@ 4.1 Experimental Setup
Implementation details and baselines. To comprehensively evaluate the effectiveness of our method, we test close-sourced LLMs GPT-3.5, GPT-4 (OpenAI et al., 2024), and open-sourced Llama2 (Touvron et al., 2023b) as the optimizer LLMs. We employ MaaO in P-tuning (Liu et al., 2023b),  Lester et al. (2021), which are the pioneering work of prompt tuning for pre-trained language models. We also apply the optimization methods to prompt tuning methods for the vision-language model, CoOp (Zhou et al., 2022b), TCP (Yao et al., 2024). CoOp is the founder of prompt tuning for vision-language models, and TCP represents one of the state-of-the-art advancements in this realm. Both methods exemplify the use of textual prompting techniques for enhancing vision-language models. For a fair comparison, we fix the original hyperparameter of previous methods, such as pre-trained backbone and prompt module design, and only apply our method as a new optimization strategy. For the configuration of Algorithm 1, the number of rounds $N$ is set as 3, and the iteration for the gradient optimizer $m$ is set as 10. All experimental results are averaged over 3 random seeds. More detailed hyperparameter settings are provided in the appendix.
Datasets. For the lauguage model, we conduct experiments over the commonly-used pre-trained model RoBERTa (Liu et al., 2019) on NLU tasks from SuperGLUE (Wang et al., 2020) to test our methods. We apply our prompt optimization algorithm to vision-language pre-trained CLIP (Radford et al., 2021) for adaptation of image classification tasks. We adopt commonly used 10 datasets to comprehensively evaluate our method, including Caltech101 (Li et al., 2004), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), FGVCAircraft (Maji et al., 2013), SUN397 (Xiao et al., 2010), UCF101 (Soomro et al., 2012), DTD (Cimpoi et al., 2014), and EuroSAT (Helber et al., 2019). Labeled few-shot samples from each class are used as training data for each dataset.
@@ 4.2 Main Results
Prompt optimization for language models. We employ the proposed optimization method for the prompt tuning of pre-trained language model RoBERTa-large (Liu et al., 2019) and evaluate on the widely used SuperGLUE (Wang et al., 2020) NLU tasks. Previous prompt tuning methods, P-tuning (Liu et al., 2023b) and  Lester et al. (2021), use only backpropagated gradient to optimize the prompt. From Table 1, our combined optimization method with GPT-4 as optimizer surpasses vanilla gradient-based optimization on four out of five tasks from SuperGLUE.
Prompt optimization for vision-language pre-trained models. We also compare with prompt tuning methods for vision-language models,  (Zhou et al., 2022b; Yao et al., 2024). From Table 2, the results on the "RN50" backbone show that our integrated optimization outperforms existing gradient-based prompt tuning methods at six out of ten benchmark datasets, and the other tasks remain close to the baseline performance. Both close-sourced GPT models and open-sourced Llama2 achieve consistent improvements, demonstrating the effectiveness of our combined optimization framework. TCP (Yao et al., 2024) is one of the state-of-the-art prompt tuning approaches with a stronger backbone. Although the absolute improvement inevitably decreases, our method still brings stable improvements on six out of ten datasets.
We also compare with methods that optimize by LLM only, e.g., Liu et al. (2023a). We list the results of the 1-shot and 4-shot settings reported in their paper since the code has not been released yet, and our reproduced results can not match those published in the paper. Although Liu et al. (2023a) achieves a completely program-free prompt learning method by LLM, their performance in the 16-shot setting is poor. This method merely relies on the inherent deductive ability of LLM, which can not make good use of information in more training samples. In our method, the gradient optimizer can promise a stable convergence by learning from more data. And instructed LLMs can exploit in a more promising sub-region of the solution space according to the intermediate results of the gradient optimizer. Thus, better performance is achieved by the collaborative optimization process.
@@
Interpretation of prompts optimized by LLM. To further analyze the contribution of LLM-based optimizer in prompt optimization, we list the prompts contained in the instruction and generated by LLM in Figure 4. In round 1, the gradient-based optimizer tend to navigate around senseless prompt tokens, e.g., "beh", "ila", etc. This phenomenon is alleviated after leveraging LLM to infer more meaningful prompts. In round 3, prompts with both interpretability and low loss values are obtained by the collaboration of gradient-based and LLM-based optimizers.
4.3 Ablation Study
Our method’s sensitivity to the choice of LLMs, prompt tuning baseline methods, and amount of training samples are already shown in Table 2. We provide more ablation results in this section. The ablation experiments are based on the CoOp baseline, using GPT-3.5 as an optimizer. More ablation studies can be found in the Appendix.
Table 3: More results demonstrating the relation of the two collaborative optimizers.
Methods EuroSAT DTD Oxford_Flowers
Single-start Gradient Optimization 69.14 54.96 85.04
Multi-start Gradient Optimization 70.64 55.08 85.45
Multi-start Gradient Optimization with Perturbations 70.33 55.42 85.71
LLM-based Optimization 49.21 44.16 67.05
Ours 71.39 56.70 86.64
More analysis of the two optimizers.
To represent the relation of the two optimizers more clearly, we provide the results in Table 3. "Single-start Gradient Optimization" refers to the basic training procedure of prompt tuning. The parameters are initialized and then trained to convergence in a single training run. To eliminate the influence of training protocol and randomness, we extend the single-start training to "Multi-start Gradient Optimization" by incorporating multiple rounds of optimization. The first round trains from initialized parameters, and the successive rounds restart training from retained parameters of the previous round. "Multi-start Gradient Optimization with Perturbations" means we add random noise values sampled from $0.01*\mathcal{N}(0,1)$ to the prompt parameters before each restart round for the opportunity to escape from local optima.
From the results, gradient-based learning itself can not significantly benefit from longer training process and random parameter perturbations. The performance gain of our method lies in the collaboration of gradient-based optimizer and the high-level guidance of LLMs.